please show all work clearly thanks
Find: 5x37x + 1 dx
37x-1 7. Find: 5 dx 8. Find: S(sin? x + cos²x) dx 9. Find: S(1 + tan’x) dx 10. Find: S(-1 - cotx) dx
Answer #1

* I-Jx²2x + 1 de +1=u Let x-4 da du I= (14-17 su du 2 ? nie for ju² (43-34² +34-1). s (u² 3 4 4 3 4 Jul de Lold 11 글이 3xl 해 U8.-S (sint cosx) da sinta + cos²x=1 I a I de We know tRER , : I - ze tc 9 . se+tank) da I J cittando de 2 2 = seixdx Ittana=s